# Thermodynamics of Kerr black hole: Tsallis-Cirto composition law and entropy quantization

### v5

@@ Abstract
The processes of splitting and merging of black holes obey the composition law generated by the Tsallis-Cirto $\delta=2$ statistics. The same composition law expresses the full entropy of the Reissner-Nordström black hole via the entropies of its outer and inner horizons. He we apply this composition law to the entropy of the Kerr black hole. As distinct from Reissner-Nordström black hole, where the full entropy depends only on mass $M$ and does not depend on its charge $Q$, the entropy of Kerr black hole is the sum of contributions from its mass $M$ and angular momentum $J$, i.e. $S(M,J)=S(M,0)+4\pi\sqrt{J(J+1)}$. Here $S(M,0)$ is the entropy of the Schwarzschild black hole. This demonstrates that when the Kerr black hole with $J\gg 1$ absorbs or emits a massless particle with spin $s_{z}=\pm 1/2$, its entropy changes by $|\Delta S|=2\pi$. We also considered the quantization of entropy suggested by the toy model, in which the black hole thermodynamics is represented by the ensemble of the Planck-scale black holes – Planckons. The Tsallis-Cirto composition law is also extended to the entropy of Kerr-Newman black hole.
@@ I Introduction
In the paper Volovik2021 , an attempt was made to consider the entropy of rotating black holes – the Kerr black hole and the Kerr-Newman black hole. The method of adiabatic transformations was used for this purpose. However, it turned out that this method is inapplicable for the rotating case. Here, we consider the entropy of rotating black holes using non-extensive Tsallis-Cirto statistics. In Sec. 2 we recall the Tsallis-Cirto type statistics, which provides a composition rule for black hole merger and splitting processes, as well as for the combination of the entropies of the inner and outer horizons of an RN black hole. This composition law is applied to Kerr and Kerr-Newman black holes in Sec. III.
@@ IV Conclusion
We applied the composition law generated by the Tsallis-Cirto $\delta=2$ statistics to the Kerr black hole and expressed the entropy of black hole via the entropies of its outer and inner horizons. In this approach we obtained that the entropy of Kerr black hole is the sum of contributions from its mass $M$ and angular momentum $J$, i.e. $S(M,J)=S(M,0)+4\pi\sqrt{J(J+1)}$, where $S(M,0)$ is the entropy of the Schwarzschild black hole. In the thermodynamic limit $J\gg 1$ this gives ${\partial S}/{\partial J}{\big|}_{M}=4\pi$. When the Kerr black hole with $J\gg 1$ absorbs or emits a massless particle with spin $s_{z}=\pm 1/2$, its entropy changes by $|\Delta S|=2\pi$.

### v6

@@ Abstract
The processes of splitting and merging of black holes obey the composition law generated by the Tsallis-Cirto $\delta=2$ statistics. The same composition law expresses the full entropy of the Reissner-Nordström black hole via the entropies of its outer and inner horizons. Here we apply this composition law to the thermodynamics of the Kerr black hole. As distinct from Reissner-Nordström black hole, where the full entropy depends only on mass $M$ and does not depend on its charge $Q$, the entropy of Kerr black hole is the sum of contributions from its mass $M$ and angular momentum $J$, i.e. $S(M,J)=S(M,0)+4\pi\sqrt{J(J+1)}$. Here $S(M,0)$ is the entropy of the Schwarzschild black hole. This demonstrates that when the Kerr black hole with $J\gg 1$ absorbs or emits a massless particle with spin $s_{z}=\pm 1/2$, its entropy changes by $|\Delta S|=2\pi$. We also considered the quantization of entropy suggested by the toy model, in which the black hole thermodynamics is represented by the ensemble of the Planck-scale black holes – Planckons. The Tsallis-Cirto composition law is also extended to the thermodynamics of Kerr-Newman black hole.
@@ I Introduction
In the paper Volovik2021 , an attempt was made to consider the entropy of rotating black holes – the Kerr black hole and the Kerr-Newman black hole. The method of adiabatic transformations was used for this purpose. However, it turned out that this method is inapplicable for the rotating case. Here, we consider the thermodynamics of rotating black holes using non-extensive Tsallis-Cirto statistics. In Sec. 2 we recall the Tsallis-Cirto type statistics, which provides a composition rule for black hole merger and splitting processes, as well as for the combination of the entropies of the inner and outer horizons of an RN black hole. This composition law is applied to Kerr and Kerr-Newman black holes in Sec. III.
@@ III.4 Entropy quantization in Planckon toy model
III.5 Thermodynamics of rotating black holes
The temperature and entropy of the RN black hole with two horizons are determined by the correlations between the inner and outer horizons. This gives rise to the thermodynamics of the RN black hole, which differs from the traditional thermodynamics determined only by the outer horizon.Davis1977 The same takes place for the Kerr black hole. Using Eq.(12) and thermodynamic relation
$dM=TdS+{\boldsymbol{\Omega}}\cdot d{\bf J}\,$, (20)
one obtains temperature of the rotating black hole:
$T=\frac{\partial M}{\partial S}\bigg|_{\bf J}=\frac{1}{8\pi GM}\,.$ (21)
Similar to the RN black hole, this temperature does not deviate from the temperature of the Schwarzschild black hole with the same mass.
Eq.(12) and thermodynamic relation Eq.(20) also determine the angular velocity of the Kerr black hole:
${\boldsymbol{\Omega}}=\frac{\partial M}{\partial{\bf J}}\bigg|_{S}=-\frac{\bf J}{2GMJ}\,.$ (22)
The angular velocity demonstrates the negative moment of inertia and it has the following connection with the temperature in Eq.(21), see also Eq.(14):
$\Omega\equiv|{\boldsymbol{\Omega}}|=4\pi T\,.$ (23)
For the Kerr-Newman black hole with entropy in Eq.(17) one has:
$T=\frac{1}{8\pi GM}\,\,,\,\,{\boldsymbol{\Omega}}=-\frac{8\pi T}{\sqrt{4J^{2}+\alpha^{2}Q^{4}}}\,{\bf J}\,.$ (24)
It should be noted that there is no discontinuity in heat capacity here, unlike the thermodynamics of rotating black hole based on the outer horizon.Davis1977
The connection between $\Omega$ and $T$ suggests the possible modification of the Zel’dovich-Starobinsky rotational superradiance.Zeldovich1971 ; Zeldovich1972 ; Starobinsky1973 ; StarobinskyChurilov1974 At external temperatures, spontaneous emission can occur even at $\omega>m\Omega$.Endlich2017 The black hole temperature in the equation (23) can play a similar role.
@@ IV Conclusion
We applied the composition law generated by the Tsallis-Cirto $\delta=2$ statistics to the Kerr black hole and expressed the entropy of black hole via the entropies of its outer and inner horizons. In this approach we obtained that the entropy of Kerr black hole is the sum of contributions from its mass $M$ and angular momentum $J$, i.e. $S(M,J)=S(M,0)+4\pi\sqrt{J(J+1)}$, where $S(M,0)$ is the entropy of the Schwarzschild black hole. In the thermodynamic limit $J\gg 1$ this gives ${\partial S}/{\partial J}{\big|}_{M}=4\pi$, while the angular velocity of the rotating black hole is expressed in terms of its temperature, $\Omega=4\pi T$. When the Kerr black hole with $J\gg 1$ absorbs or emits a massless particle with spin $s_{z}=\pm 1/2$, its entropy changes by $|\Delta S|=2\pi$.